If I make $300 weekly, and have a 6% pay raise plus $12 for traveling costs, how much of an increase will I receive?
Answer by stanbon(75887)

If I make $300 weekly, and have a 6% pay raise plus $12 for traveling costs, how much of an increase will I receive?
-------------------
Increase = 0.06*300 + 12 = 18 + 12 = $30
---------------------------------------
Cheers,
Stan H.
-------------